OpenStudy (anonymous):

Which relationship is always correct for the angles x, y, and z of triangle ABC?

OpenStudy (anonymous):

x + z = y y + z = x x + y + z = 180 degrees x + y + z = 90 degrees

OpenStudy (anonymous):

angle at A is 180-(x) angle at A is 180-(y+z) so x=y+z